# FusionPortableV2: A Unified Multi-Sensor Dataset for Generalized SLAM Across Diverse Platforms and Scalable Environments

### v1

@@ 7.3.1 Data Preparation:
Each frame image is accompanied by a GT depth image of identical size for evaluation. Depth images are produced by projecting point clouds, generated by FAST-LIO2 (Xu et al., 2022) through IMU interpolastion, onto these frames:
$D_{gt}(\mathbf{x})=Z,\ \ \ \mathbf{x}=\lfloor\pi(\mathbf{p}^{c})\rceil,\ \ \ \mathbf{p}^{c}=\mathbf{R}^{c}_{l}\mathbf{p}^{l}+\mathbf{t}^{c}_{l},$ (4)
@@ 7.3.2 Experiment Setting:
Monocular depth estimation tests are essential for evaluating a system’s ability to perceive relative object distances from a single camera view, a key aspect of understanding spatial relationships. Based on the FSNet baseline depth estimation modal (Liu et al., 2023), we fine-tune the modal with our dataset. We organize the train-validation data into two groups. The first group allocates $70\%$ of handheld indoor sequence data (i.e., handheld_room00 and handheld_escalator00) for training, and combines $30\%$ of this data with all vehicle-related outdoor sequences for validation. The second group uses $70\%$ of vehicle sequence data (i.e., vehicle_campus00 and vehicle_parking00) for training, and blends $30\%$ of this data with all handheld sequences for validation. We train the FSNet with these groups of data respectively and obtain two modesls: FSNet-Handheld and FSNet-Vehicle.
@@ 7.3.3 Evaluation:
We assess models’ performance of unsupervised monocular depth prediction models use the proposed scale-invariant metrics in (Eigen et al., 2014): Absolute Relative Difference (ARD), Squared Relative Difference (SRD), Root Mean Squared Error (RMSE)-linear, RMSE-log, and Threshold. These metrics are defined as follow:
@@ 7.4.1 Calibration:
Achieving the life-long sensor calibration poses significant challenges (Maddern et al., 2017). We try our best to provide the best estimate of calibration parameters. Calibration was performed each time when the data collection platform changed, employing SOTA methods for parameter adjustments, which were also manually verified and fine-tuned. For extrinsic parameters difficult to estimate, such as the relative transformation between specific components, we refer to the CAD model. Efforts were made to reinforce the mechanical structure and minimize external disturbances during the data collection process. Nevertheless, it is acknowledged that high accuracy for specific traversals cannot be assured. Users are encouraged to use our calibration estimates as initial values and explore innovative approaches for long-term extrinsic calibration. To aid in these endeavors, we provide raw calibration data and reports, allowing users to develop their methodologies and consider our estimates as a foundational benchmark.
@@ 7.4.3 Partial Loss of Information:
In the construction of the dataset, real-world challenges have led to partial loss of sensor information in certain sequences, reflecting practical issues encountered during robotic deployment. Specifically, for two sequences captured using the Unmanned Ground Vehicle (UGV), the wheel encoder driver was not activated correctly, resulting in the absence of wheel encoder data. Additionally, one sequence from the legged robot experienced brief interruptions in data transmission, amounting to several seconds of lost data, due to a faulty RJ45 network port connection. These instances underscore the importance of robustness in algorithm development to handle incomplete or missing sensor data in realistic operational conditions.
@@ 7.4.4 Camera Exposure Setting:
7.4.5 Scale Discrepancy:
We recognize that the notable limitation of the proposed dataset is the scale, as compared with vast datasets that are trained by exiting foundation models (Brohan et al., 2023). This disparity in scale poses challenges for training robust, generalized SLAM models. Nonetheless, this paper marks a significant initial endeavor in addressing basic challenges in data collection for field robots, including system integration and data postprocessing. This will lay the groundwork for future explorations and expansions in this field.
@@ 8 Conclusion and Future Work
In this paper, we present the FusionPortableV2 dataset, a comprehensive multi-sensor collection designed to advance research in SLAM and mobile robot navigation. The dataset is built around a compact, multi-sensor device that integrates IMUs, stereo cameras (both frame- and event-based), LiDAR, and GNSS, all of which are carefully calibrated and synchronized. This primary device is deployed on various platforms, including legged robot, low-speed UGV, and high-speed vehicle, each equipped with additional platform-specific sensors such as wheel encoders and legged sensors. Furthermore, the FusionPortableV2 dataset lies in its diverse range of environments, spanning indoor spaces, grasslands, campuses, parking lots, tunnels, downhill roads, and highways. This environmental diversity allows the dataset to challenge existing SLAM and navigation technologies with realistic scenarios that involve dynamic objects and variable lighting conditions. To ensure the dataset’s utility for the research community, we have meticulously designed $27$ sequences, totaling $2.5$ hours of data, and have provided GT data for objective evaluation of SOTA methods in localization, mapping, and monocular depth estimation.
As we explore future directions for this work, we aim to enhance this dataset’s applicability beyond SLAM, by developing novel navigation methods based on the proposed dataset. We continue to improve the quality of the data and integration of the system for easier manufacture by nonexpert users. Alongside this dataset, we release our implementation details and tools to encourage further research advancements.
@@
This research greatly benefited from the guidance and expertise of many contributors. We extend our profound gratitude to colleagues: Ruoyu Geng, Lu Gan, Bowen Yang, Tianshuai Hu, Mingkai Jia, Mingkai Tang, Yuanhang Li, Shuyang Zhang, Bonan Liu, Jinhao He, Ren Xin, Yingbing Chen, etc. from HKUST and HKUSTGZ for significantly improving our dataset’s quality and applicability, especially in data collection. Special acknowledgment goes to the BIM Lab at HKUST, particularly Prof. Jack Chin Pang Cheng and his students, for their crucial BLK360 scanning expertise and insights that were essential in shaping our dataset. The authors also gratefully acknowledge Dr. Thien-Minh Nguyen (NTU) and Ms. Qingwen Zhang (KTH) for their insightful feedback and technical support; Prof. Dimitrios Kanoulas (UCL), Prof. Martin Magnusson (Örebro University), Prof. Hong Zhang (Southern University of Science and Technology), and Prof. Peng Yin (CityU Hong Kong) for their invaluable dataset writing advice; Mr. Seth G. Isaacson (University of Michigan) for his suggestions on dataset structure. The author(s) received no financial support for the research, authorship, and/or publication of this article.

### v2

@@ 7.3.1 Data Preparation:
Each frame image is accompanied by a GT depth image of identical size for evaluation. Depth images are produced by projecting point clouds, generated by FAST-LIO2 (Xu et al., 2022) through IMU interpolation, onto these frames:
$D_{gt}(\mathbf{x})=Z,\ \ \ \mathbf{x}=\lfloor\pi(\mathbf{p}^{c})\rceil,\ \ \ \mathbf{p}^{c}=\mathbf{R}^{c}_{l}\mathbf{p}^{l}+\mathbf{t}^{c}_{l}.$ (4)
@@ 7.3.2 Experiment Setting:
Monocular depth estimation tests are essential for evaluating a system’s ability to perceive relative object distances from a single camera view, a key aspect of understanding spatial relationships. Based on the FSNet that is a self-supervised depth estimation model (Liu et al., 2023), we fine-tune this model with our dataset using GT poses. We organize the train-validation data into two groups. The first group allocates $70\%$ of handheld indoor sequence data (i.e., handheld_room00 and handheld_escalator00) for training, and combines $30\%$ of this data with all vehicle-related outdoor sequences for validation. The second group uses $70\%$ of vehicle sequence data (i.e., vehicle_campus00 and vehicle_parking00) for training, and blends $30\%$ of this data with all handheld sequences for validation. We train the FSNet with these groups of data respectively and obtain two models: FSNet-Handheld and FSNet-Vehicle.
@@ 7.3.3 Evaluation:
We assess models’ performance of unsupervised monocular depth prediction models use the proposed scale-invariant metrics in (Eigen et al., 2014): Absolute Relative Difference (ARD), Squared Relative Difference (SRD), Root Mean Squared Error (RMSE)-linear, RMSE-log, and Threshold. They are defined as follows:
@@ 7.4.1 Calibration:
Achieving the life-long sensor calibration poses significant challenges (Maddern et al., 2017). We try our best to provide the best estimate of calibration parameters. Calibration was performed each time when the data collection platform changed, employing SOTA methods for parameter adjustments, which were also manually verified and fine-tuned. For extrinsic parameters difficult to estimate, such as the relative transformation between specific components, we refer to the CAD model. Efforts were made to reinforce the mechanical structure and minimize external disturbances during the data collection process. Nevertheless, it is acknowledged that high accuracy for specific traversals cannot be assured. Users are encouraged to use our calibration estimates as initial values and explore innovative approaches for long-term extrinsic calibration, such as (Wu et al., 2021; Ulrich and Hillemann, 2023; Luo et al., 2024). To aid in these endeavors, we provide raw calibration data and reports, allowing users to develop their methodologies and consider our estimates as a foundational benchmark.
@@ 7.4.3 Partial Loss of Information:
In the construction of the dataset, real-world challenges have led to partial loss of sensor information in certain sequences, reflecting practical issues encountered during robotic deployment. Specifically, in ugv_transition00 and ugv_transition01, the wheel encoder driver was not activated correctly, resulting in the absence of wheel encoder data. Additionally, the legged_grass00 sequence experienced a brief interruption in data transmission, amounting to several seconds of lost data, due to a loose RJ45 network port connection. These instances underscore the importance of robustness in algorithm development to handle incomplete or missing sensor data in realistic operational conditions.
@@ 7.4.4 Camera Exposure Setting:
7.4.5 Limited Diversity and Volume:
Drawing inspiration from foundational models in computer vision and natural language processing, developing a robot-agnostic general model for robotics could be an exciting avenue for future research (Padalkar et al., 2023; Khazatsky et al., 2024; Shah et al., 2023a). However, acquiring a large, diverse real-world dataset poses significant challenges. We address key problems in data collection for field robots, including system integration and data postprocessing. But we acknowledge the limitations in our dataset’s scale and diversity: (1) Platforms: Our dataset does not include drones, underwater robots, or multi-robot systems. (2) Scenarios: Critical environments such as factories, forests, and underground mines, which are also important for SLAM but not presented. (3) Time Periods: The dataset was collected over a short time frame, lacking day-night cycles, seasonal variations, and structural changes, which are crucial for place recognition and long-term SLAM (Carlevaris-Bianco et al., 2016; Barnes et al., 2020). (4) Volume: The dataset is relatively small, with a limited number of sequences and short durations. FusionPortableV2 is just the beginning. As more researchers and institutions contribute to this field, we anticipate the development of larger, more diverse datasets that will support the creation of a truly generalizable robotic model.
@@ 8 Conclusion and Future Work
This paper presents the FusionPortableV2 dataset, a comprehensive multi-sensor collection designed to advance research in SLAM and mobile robot navigation. The dataset is built around a compact, multi-sensor device that integrates IMUs, stereo cameras (both frame-based and event-based), LiDAR, and INS, all carefully calibrated and synchronized. This primary device is deployed on various platforms, including a legged robot, a low-speed UGV, and a high-speed vehicle, each equipped with additional platform-specific sensors such as wheel encoders and legged sensors. The FusionPortableV2 dataset features a diverse range of environments, spanning indoor spaces, grasslands, campuses, parking lots, tunnels, downhill roads, and highways. This environmental diversity challenges existing SLAM and navigation technologies with realistic scenarios involving dynamic objects and variable lighting conditions. To ensure the dataset’s utility for the research community, we have meticulously designed 27 sequences, totaling 2.5 hours of data, and provided ground truth data for the objective evaluation of SOTA methods in localization, mapping, and monocular depth estimation.
As we explore future directions, we aim to enhance this dataset’s applicability beyond SLAM by developing novel navigation methods based on the proposed dataset. We will continue to improve the quality of the data and the integration of the system to facilitate easier use by non-expert users. Alongside this dataset, we also release our implementation details and tools to encourage further research advancements.
@@
This research greatly benefited from the guidance and expertise of many contributors. We extend our profound gratitude to colleagues: Ruoyu Geng, Lu Gan, Bowen Yang, Tianshuai Hu, Mingkai Jia, Mingkai Tang, Yuanhang Li, Shuyang Zhang, Bonan Liu, Jinhao He, Ren Xin, Yingbing Chen, etc. from HKUST and HKUSTGZ for their suggestions in improving our dataset’s quality. Special acknowledgment goes to the BIM Lab at HKUST, particularly Prof. Jack Chin Pang Cheng and his students, for their crucial BLK360 scanning expertise and insights that were essential in shaping our dataset. The authors also gratefully acknowledge Dr. Thien-Minh Nguyen (NTU) and Ms. Qingwen Zhang (KTH) for their insightful feedback and technical support; Prof. Dimitrios Kanoulas (UCL), Prof. Martin Magnusson (Örebro University), Prof. Hong Zhang (Southern University of Science and Technology), and Prof. Peng Yin (CityU Hong Kong) for their invaluable dataset writing advice; Mr. Seth G. Isaacson (University of Michigan) for his suggestions on dataset structure. ; and OpenAI’s GPT for enhancing the paper’s linguistic quality. The author(s) received no financial support for the research, authorship, and/or publication of this article.